OpenStudy (anonymous):

sum of (n)/(e^(3n)), n=1 to inf

OpenStudy (anonymous):

PS: Use the integral test to see if the series converges or diverges.

OpenStudy (anonymous):

Σ(n = 1 to ∞) (n+1)(2n+1)/(n+2)! = Σ(n = 1 to ∞) ((n+2) - 1)(2(n+2) - 3)/(n+2)! = Σ(n = 1 to ∞) (2(n+2)^2 - 5(n+2) + 3)/(n+2)! = Σ(n = 1 to ∞) [2(n+2)/(n+1)! - 5/(n+1)! + 3/(n+2)!] = Σ(n = 1 to ∞) [2((n+1)+1)/(n+1)! - 5/(n+1)! + 3/(n+2)!] = Σ(n = 1 to ∞) [2/n! + 2/(n+1)! - 5/(n+1)! + 3/(n+2)!] = 2 * Σ(n = 1 to ∞) 1/n! + Σ(n = 1 to ∞) [-3/(n+1)! + 3/(n+2)!] = 2(e - 1/0!) + Σ(n = 1 to ∞) [-3/(n+1)! + 3/(n+2)!], since Σ(n = 0 to ∞) 1/n! = e = (2e - 2) + lim(k→∞) [(-3/2! + 3/3!) + (-3/3! + 3/4!) + ... + (-3/(k+1)! + 3/(k+2)!)] = (2e - 2) + lim(k→∞) (-3/2! + 3/(k+2)!), since all other terms cancel in pairs = (2e - 2) + (-3/2! + 0) = 2e - 7/2. -------------- I hope this helps! 2 Comment . Other Answers (1) ? answered 9 months ago your question is not clear 1 Comment . Sign In to add your answer . Ask a Question Related Questions Series Sum(infinity; n=1) series 1/n(n+1)(n+2) converges or diverges? Sum from 1 to infinity of 1/(n^2+1)? Find the sum of the series sum (2^(n-1))/(5^(n+1)) n=1 to infinity? Sum of convergent series: ∑ (n=2 to infinity) 1 / (n^2 -1)? Here's a fun 1! test this series for convergence sum from n=1 to infinity of (2^1/n - 1)? Discover Questions Find an element of the largest order in S_n for n = 1,2,3...10? Help with area in polar coordinates problem? Can someone help with a little algebra? What is after 999 Novengoogolplexion? Terms Privacy AdChoices RSS Σ(n = 1 to ∞) (n+1)(2n+1)/(n+2)! = Σ(n = 1 to ∞) ((n+2) - 1)(2(n+2) - 3)/(n+2)! = Σ(n = 1 to ∞) (2(n+2)^2 - 5(n+2) + 3)/(n+2)! = Σ(n = 1 to ∞) [2(n+2)/(n+1)! - 5/(n+1)! + 3/(n+2)!] = Σ(n = 1 to ∞) [2((n+1)+1)/(n+1)! - 5/(n+1)! + 3/(n+2)!] = Σ(n = 1 to ∞) [2/n! + 2/(n+1)! - 5/(n+1)! + 3/(n+2)!] = 2 * Σ(n = 1 to ∞) 1/n! + Σ(n = 1 to ∞) [-3/(n+1)! + 3/(n+2)!] = 2(e - 1/0!) + Σ(n = 1 to ∞) [-3/(n+1)! + 3/(n+2)!], since Σ(n = 0 to ∞) 1/n! = e = (2e - 2) + lim(k→∞) [(-3/2! + 3/3!) + (-3/3! + 3/4!) + ... + (-3/(k+1)! + 3/(k+2)!)] = (2e - 2) + lim(k→∞) (-3/2! + 3/(k+2)!), since all other terms cancel in pairs = (2e - 2) + (-3/2! + 0) = 2e - 7/2. -------------- I hope this helps! 2 Comment . Other Answers (1) ? answered 9 months ago your question is not clear 1 Comment . Sign In to add your answer . Ask a Question Related Questions Series Sum(infinity; n=1) series 1/n(n+1)(n+2) converges or diverges? Sum from 1 to infinity of 1/(n^2+1)? Find the sum of the series sum (2^(n-1))/(5^(n+1)) n=1 to infinity? Sum of convergent series: ∑ (n=2 to infinity) 1 / (n^2 -1)? Here's a fun 1! test this series for convergence sum from n=1 to infinity of (2^1/n - 1)? Discover Questions Find an element of the largest order in S_n for n = 1,2,3...10? Help with area in polar coordinates problem? Can someone help with a little algebra? What is after 999 Novengoogolplexion? Terms Privacy AdChoices RSS